Calculate the formula mass of caso4. what is the molarity of caso4 in a solution containing 1.2 gram of cso4 in a volume of 50 ml how many grams of caso4 are in 50 ml of 0.086 m caso4?


1
Expert's answer
2022-04-16T04:09:04-0400

Mr(CaSO4) = Ar(Ca) + Ar(S) + 4*Ar(O) = 40 + 32 + 4*16 = 136 amu


n(CaSO4) = m(CaSO4)/Mr(CaSO4) = 1.2/136 = 0.008826 mol

50 mL = 0.05 L

c(CaSO4) = n(CaSO4)/V(solution) = 0.008826/0.05 = 0.1765 M


n(CaSO4) = c(CaSO4)*V(solution) = 0.086*0.05 = 0.0043 mol

m(CaSO4) = n(CaSO4)*Mr(CaSO4) = 0.0043*136 = 0.5848 g